Find the Laplace transform of the piecewise continuous function
f(x) = {1, 0 ≤ t < 1}
{-3e^-t, t ≥ 1}
1
Expert's answer
2020-07-26T17:45:39-0400

"F(s)=\int\limits_0^\infty e^{-t\cdot s}\cdot f(t)\,dt=\\

=\int\limits_0^1 e^{-t\cdot s}\cdot f(t)\,dt +\int\limits_1^\infty e^{-t\cdot s}\cdot f(t)\,dt=\\
=\int\limits_0^1 e^{-t\cdot s}\cdot 1\cdot\,dt +\int\limits_1^\infty e^{-t\cdot s}\cdot (-3\cdot e^{-t})\,dt=\\
=\frac{1-e^{-s}}{s}-\frac{3\cdot e^{-s-1}}{s+1} ;"